# RGNMR: A Gauss-Newton method for robust matrix completion with theoretical guarantees

### v1

@@ 3 Recovery guarantees
  • $\mu$ incoherence parameter of $L^{*}$
  • $\alpha$ a bound on the fraction of outliers in each row/column
  • $\gamma$ over-removal factor
  • $\delta$ neighborhood parameter
@@
Table 1: Recovery guarantees requirements of various RMC methods, up to multiplicative constant factors. The weakest conditions for each type of methods are in bold.
@@
In each row and column, the fraction of observed entries which are corrupted is bounded. Formally, for a known $10<\alpha<$ 1, we assume that $S^{*}\in\mathcal{S}_{\alpha}^{\Omega}$, where
@@
Assumption 1 was introduced by Candes and Recht, (2012). Assumptions 2 and 3 or variants thereof are common in theoretical analyses of matrix completion, see for example (Candès et al.,, 2011; Yi et al.,, 2016; Cherapanamjeri et al.,, 2017; Wang and Wei,, 2024).
Similar to other works on robust matrix completion, we derive guarantees for a slightly modified version of our proposed method. Specifically, we make the following two modifications. First, based on assumption 3 that the fraction of corrupted entries in each row and column is sufficiently small, we estimate the set of corrupted entries $\Lambda_{*}$ in a different way. To this end we introduce the following thresholding operator $\mathcal{T}_{\theta}(A,\Omega)$. Given an input matrix $A$, a subset $\Omega$, and a parameter $\theta\in(0,1)$, the operator $\mathcal{T}_{\theta}(A,\Omega)$ keeps only those entries $(i,j)\in\Omega$ which belong to the largest $\theta$-fraction of entries in both the respective row $i$ and column $j$, and zeros out all the remaining entries. In case of entries having identical magnitude, ties are broken arbitrarily. Formally,
@@
For RGNMR to be able to remove all outliers, a necessary condition is that $\theta\geq\alpha$. In what follows, we choose $\theta=\gamma\alpha$ where $\gamma\geq 1$ is the over-removal factor.
@@
$\mathcal{B}_{\mu}$ =$=\left\{\begin{pmatrix}U\\ V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,\|U\|_{2,\infty}\leq\sqrt{\frac{3\mu r\sigma_{1}^{*}}{n_{1}}},\quad\|V\|_{2,\infty}\leq\sqrt{\frac{3\mu r\sigma_{1}^{*}}{n_{2}}}\right\},$ (9)
where the constant $3$ is arbitrary. Instead of eq. 1, for a suitably chosen $\delta>0$, the modified RGNMR updates the factor matrices as follows,
@@
Similar constraints were employed by Zilber and Nadler, (2022) and Keshavan et al., (2010), in deriving theoretical guarantees for their (non-robust) matrix completion algorithms. As these constraints are quadratic, the above problem may be written as a regularized least squares problem with quadratic regularization terms. Hence, it can be solved computationally efficiently. The modified RGNMR is described in Algorithm 2.
@@
$\mathcal{B}_{\text{err}}(\epsilon)$ =$=\left\{\begin{pmatrix}U,V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,\|UV^{\top}-L^{*}\|_{F}\leq\epsilon\sigma_{r}^{*}\right\}.$ (11)
@@
=$\mathcal{B}^{*}=\left\{\begin{pmatrix}U,V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,UV^{\top}=L^{*}\right\}.$ (12)
@@
$\mathcal{B}_{\textnormal{bln}}(\delta)$ =$=\left\{\begin{pmatrix}U,V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,\|U^{\top}U-V^{\top}V\|_{F}\leq\delta\sigma_{r}^{*}\right\}.$ (13)
@@
Theorem 3.1 holds for $\frac{25\sigma_{r}^{*}}{c_{e}^{2}\kappa}\leq\delta\leq\frac{\sigma_{r}^{*}}{c_{e}\kappa}$. In principle, to provide a valid value of $\delta$ in this interval requires knowledge of $\sigma_{1}^{*}$ and $\sigma_{r}^{*}$. We note that several other works also required knowledge of $\sigma_{1}^{*}$ or $\sigma_{r}^{*}$ or both for their methods, see (Yi et al.,, 2016; Sun and Luo,, 2016; Cherapanamjeri et al.,, 2017). In fact, for our approach it suffices to estimate these two singular values up to constant multiplicative factors. Moreover, within the proof of Theorem 3.2, we show that under its assumptions, it is indeed possible to estimate $\sigma_{1}^{*}$ and $\sigma_{r}^{*}$ up to constant factors.
@@
Theorems 3.1 and 3.2 require that the parameters $\mu$ and $\alpha$ are known. Similar assumptions were made in previous works, e.g. (Yi et al.,, 2016; Cherapanamjeri et al.,, 2017; Sun and Luo,, 2016; Zilber and Nadler,, 2022).
@@
For the vanilla matrix completion problem, where the observed entries are not corrupted, the smallest sample complexity requirement for factorization based methods was derived by Zheng and Lafferty, (2016). Remarkably, Theorem 3.1 matches this result for RGNMR even in the presence of corrupted entries. The number of samples required by Theorem 3.1 is smaller than those required for other RMC methods (Yi et al.,, 2016; Wang and Wei,, 2024) and if $\kappa<C\log(\mu r\sigma_{1}^{*})\log n_{1}$, for a suitable constant $C$, it is also smaller than the requirements in (Cherapanamjeri et al.,, 2017).
@@
In terms of the fraction $\alpha$ of corrupted entries, the recovery guarantee with the highest $\alpha$ for factorization based methods was derived by Tong et al., (2021) and by Cai et al., (2024), under the assumption that the matrix is fully observed. Theorem 3.1 improves upon this result by a factor of $\mathcal{O}(\sqrt{r})$, even though the matrix is only partially observed. Hence our recovery guarantee holds under the weakest known conditions for factorization based methods. For general RMC methods the recovery guarantee with the highest $\alpha$ was derived by Cherapanamjeri et al., (2017). Theorem 3.1 requires a smaller $\alpha$ than in (Cherapanamjeri et al.,, 2017) by a factor of $\mathcal{O}(\kappa)$. Table 1 summarizes the sample complexity and corruption rate requirements of different methods.
@@
Next, for constructing a suitable initialization the requirements on $\alpha$ and $p$ in Theorem 3.2 are more stringent than those for the recovery guarantee of Theorem 3.1. This separation is common in other works on matrix completion (Zheng and Lafferty,, 2016; Yi et al.,, 2016; Zilber and Nadler,, 2022). In terms of sample complexity Theorem 3.2 still improves upon the results in (Yi et al.,, 2016; Wang and Wei,, 2024) and in some parameter regimes also upon the result in (Cherapanamjeri et al.,, 2017). In terms of the fraction of outliers the condition in Theorem 3.2 is more stringent than in (Yi et al.,, 2016) and (Cherapanamjeri et al.,, 2017). Since empirically RGNMR succeeds even with random initialization we believe the condition imposed by Theorem 3.1 can be relaxed. This would allow for less stringent conditions in Theorem 3.2. We leave this for future work.
Limitations and future research.
@@
Although our theoretical analysis improves upon existing results, it does not fully account for the remarkable empirical performance of RGNMR. For example, the conditions in Theorem 3.1 on both $\alpha$ and $p$ depend on the condition number $\kappa$ while empirically the performance of RGNMR is not affected by $\kappa$. In addition, our analysis did not consider the case of overparameterization. These remains open directions for future theoretical analysis of RGNMR.
@@
4 Numerical results
This section details the simulations conducted to compare the various RMC methods. Their results are illustrated in Figures 1, 2, 3 and in Appendix B. Each instance consists of generating a random matrix $L^{*}\in\mathbb{R}^{n_{1}\times n_{2}}$ of a given rank $r$, condition number $\kappa$, subset of observed entries $\Omega\subseteq[n_{1}]\times[n_{2}]$ and a corruption matrix $S^{*}$. We generate $L^{*}$ following (Tong et al.,, 2021). Specifically, we construct $U\in\mathbb{R}^{n_{1}\times r},V\in\mathbb{R}^{n_{2}\times r}$ with entries $i.i.d.$ from the standard normal distribution and orthonormalize their columns. We then construct $\Sigma\in\mathbb{R}^{r\times r}$ as a diagonal matrix with $r$ evenly spaced values between $1$ and $\frac{1}{\kappa}$ on the diagonal. We set $L^{*}=U\Sigma V^{\top}$. Next, to generate $\Omega$ we follow (Zilber and Nadler,, 2022). Given an oversampling ratio $\rho$ we sample $\Omega$ of size $\lfloor\rho r(n_{1}+n_{2}-r)\rfloor$ randomly without replacement and verify that it contains at least $r$ entries in each column and row. To construct the corruption matrix $S^{*}$ we first sample its support $\Lambda_{*}\subseteq\Omega$ as follows. Given a fraction $\alpha$ of corrupted entries we sample $\lfloor\alpha\cdot|\Omega|\rfloor$ entries out of $\Omega$ randomly without replacement. We then verify that the set $\Omega\setminus\Lambda_{*}$ contains at least $r$ entries in each column and row. Similar to (Cai et al.,, 2024; Wang and Wei,, 2024), we sample the value of each corruption entry uniformly between $-\max_{i,j}|L^{*}_{i,j}|$ and $\max_{i,j}|L^{*}_{i,j}|$. Each method is then provided with the matrix $\mathcal{P}_{\Omega}(L^{*}+S^{*})$, the set of entries $\Omega$ and the matrix rank $r$.
We compare RGNMR to AOP (Yan et al.,, 2013), RMC (Cambier and Absil,, 2016), RPCA-GD (Yi et al.,, 2016), RRMC (Cherapanamjeri et al.,, 2017), HUB (Ruppel et al.,, 2020) and HOAT (Wang et al.,, 2023). In addition, as described in Section 2.1, we compare RGNMR to RGNMR-BS. Implementations of all methods except RRMC are publicly available. An implementation of RRMC was kindly provided to us by Wang and Wei, (2024). MATLAB and Python implementations of RGNMR and RGNMR-BS are given in the supplementary and will be made publicly available upon publication. All simulations were run on a 2.1GHz Intel Core i7 CPU with 32GB of memory.
Each algorithm was executed with its default parameters, with the following exceptions: (i) To improve RPCA-GD and HUB performance we increased their number of iterations by using stricter convergence threshold than their default; (ii) In RPCA-GD we employed an over removal factor $\gamma$ of $4$ as it significantly improved its performance in our simulations; (iii) We tuned the parameter $\xi$ in HOAT to $30$; (iv) We provided AOP the exact number of corrupted entries; (v) For RRMC we use the same parameters used in (Wang and Wei,, 2024).
In Appendix B we illustrate the performance of RGNMR under a broad set of additional scenarios. Specifically, Figure 5 shows that RGNMR can handle a large fraction of corrupted entries. Figure 6 shows that RGNMR outperforms other methods also under non-uniform sampling. Figure 7 presents its stability under additive noise and Figure 8 demonstrates that RGNMR can recover matrices with a high rank even if the number of observed entries is relatively small. Figure 9 illustrates that RGNMR performs well on a real dataset involving background extraction in a video. This is a standard benchmark for RMC methods (Yi et al.,, 2016; Cherapanamjeri et al.,, 2017; Huang et al.,, 2021; Cai et al.,, 2024). Finally, Figure 10 show the runtime of RGNMR as a function of matrix size. As the figure demonstrates, for an $n\times n$ matrix, the runtime of RGNMR grows quadratically with $n$.
@@ Appendix A Initialization procedure
  • $\{X_{i,j}\mid(i,j)\in\Omega\}$ observed entries
@@
  • $r$ rank of $L^{*}$
  • $p$ probability of observing an entry
  • $\alpha$ the maximal fraction of outliers in each row/column
@@
Theorem 3.1 shows that starting from a sufficiently accurate balanced initialization with bounded row norms RGNMR recovers $L^{*}$. Therefore, the goal of our initialization procedure is to construct such initialization. Our proposed scheme starts by constructing an initial estimator of $S^{*}$, which we denote by $S_{init}$. To this end, the function $\mathcal{T}_{\alpha}$ (defined in eq. 6) is applied to $\mathcal{P}_{\Omega}(X)$. Formally,
@@
[…]

### v2

@@ 3 Recovery guarantees
  • $\mu$ - incoherence parameter of $L^{*}$
  • $\alpha$ - a bound on the fraction of outliers in each row/column
  • $\gamma$ - over-removal factor
  • $\delta$ - neighborhood parameter
@@
Table 1: Recovery guarantees requirements of various RMC methods, up to multiplicative constant factors. The weakest conditions for each category of methods are in bold.
@@
In each row and column, the fraction of observed entries which are corrupted is bounded. Formally, for a known $0<\alpha<1$, we assume that $S^{*}\in\mathcal{S}_{\alpha}^{\Omega}$, where
@@
Assumption 1 was introduced by Candes and Recht, (2012). Assumptions 2 and 3 or variants thereof are common in theoretical analyses of matrix completion (see for example, Candès et al.,, 2011; Yi et al.,, 2016; Cherapanamjeri et al.,, 2017; Wang and Wei,, 2024).
Similar to other works on robust matrix completion, we derive guarantees for a slightly modified variant of our proposed method. Specifically, we make the following two modifications.
First, the set of corrupted entries $\Lambda_{*}$ is estimated differently. In the original method $\Lambda_{*}$ is estimated as the set of $k$ largest residual entries. However, motivated by Assumption 3 that the fraction of corrupted entries in each row and column is bounded, in the modified algorithm we remove only a bounded number of entries from each row and column. To this end we introduce the following thresholding operator $\mathcal{T}_{\theta}(A,\Omega)$. Given an input matrix $A$, a subset $\Omega$, and a parameter $\theta\in(0,1)$, the operator $\mathcal{T}_{\theta}(A,\Omega)$ kepdf only those entries $(i,j)\in\Omega$ which belong to the largest $\theta$-fraction of entries in both the respective row $i$ and column $j$, and zeros out all the remaining entries. In case of entries having identical magnitude, ties are broken arbitrarily. Formally,
@@
As in the original method, entries are still removed based on the magnitude of their residual but in this way the number of removed entries from each row and column is bounded. For RGNMR to be able to remove all outliers, a necessary condition is that $\theta\geq\alpha$. In what follows, we choose $\theta=\gamma\alpha$ where $\gamma\geq 1$ is the over-removal factor.
@@
$\mathcal{B}_{\mu}$ $=\left\{\begin{pmatrix}U\\ V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,\|U\|_{2,\infty}\leq\sqrt{\frac{3\mu r\sigma_{1}^{*}}{n_{1}}},\quad\|V\|_{2,\infty}\leq\sqrt{\frac{3\mu r\sigma_{1}^{*}}{n_{2}}}\right\},$ (9)
where the constant $3$ is arbitrary. Instead of Eq. (1), for a suitably chosen $\delta>0$, the modified RGNMR updates the factor matrices as follows,
@@
Similar constraints were employed by Zilber and Nadler, (2022) and Keshavan et al., (2010), in deriving theoretical guarantees for their (non-robust) matrix completion algorithms. As these constraints are quadratic, the above problem may be written as a convex optimization problem with quadratic regularization terms. Hence, it can be solved computationally efficiently. The modified RGNMR is described in Algorithm 2.
@@
$\mathcal{B}_{\text{err}}(\epsilon)$ $=\left\{\begin{pmatrix}U,V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,\|UV^{\top}-L^{*}\|_{F}\leq\epsilon\sigma_{r}^{*}\right\}.$ (11)
@@
$\mathcal{B}^{*}=\left\{\begin{pmatrix}U,V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,UV^{\top}=L^{*}\right\}.$ (12)
@@
$\mathcal{B}_{\textnormal{bln}}(\delta)$ $=\left\{\begin{pmatrix}U,V\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}\,\middle|\,\|U^{\top}U-V^{\top}V\|_{F}\leq\delta\sigma_{r}^{*}\right\}.$ (13)
@@
Theorem 3.1 holds for $\frac{25\sigma_{r}^{*}}{c_{e}^{2}\kappa}\leq\delta\leq\frac{\sigma_{r}^{*}}{c_{e}\kappa}$. In principle, to provide a valid value of $\delta$ in this interval requires knowledge of $\sigma_{1}^{*}$ and $\sigma_{r}^{*}$. We note that several other works also required knowledge of $\sigma_{1}^{*}$ or $\sigma_{r}^{*}$ or both for their methods (see Yi et al.,, 2016; Sun and Luo,, 2016; Cherapanamjeri et al.,, 2017). In fact, for our approach it suffices to estimate these two singular values up to constant multiplicative factors. Moreover, within the proof of Theorem 3.2, we show that under its assumptions, it is indeed possible to estimate $\sigma_{1}^{*}$ and $\sigma_{r}^{*}$ up to constant factors.
@@
Theorems 3.1 and 3.2 require that the parameters $\mu$ and $\alpha$ are known. Similar assumptions were made in previous works (e.g. Yi et al.,, 2016; Cherapanamjeri et al.,, 2017; Sun and Luo,, 2016; Zilber and Nadler,, 2022).
@@
For the vanilla matrix completion problem, where the observed entries are not corrupted, the smallest sample complexity requirement for factorization based methods was derived by Zheng and Lafferty, (2016). Remarkably, Theorem 3.1 matches this result for RGNMR even in the presence of corrupted entries. The number of samples required by Theorem 3.1 is smaller than those required for other RMC methods (Yi et al.,, 2016; Wang and Wei,, 2024) and if $\kappa<C\log(\mu r\sigma_{1}^{*})\log n_{1}$, for a suitable constant $C$, it is also smaller than the requirements in Cherapanamjeri et al., (2017).
@@
In terms of the fraction $\alpha$ of corrupted entries, the recovery guarantee with the highest $\alpha$ for factorization based methods was derived by Tong et al., (2021) and by Cai et al., (2024), under the assumption that the matrix is fully observed. Theorem 3.1 improves upon this result by a factor of $\mathcal{O}(\sqrt{r})$, even though the matrix is only partially observed. Hence our recovery guarantee holds under the weakest known conditions for factorization based methods. For general RMC methods the recovery guarantee with the highest $\alpha$ was derived by Cherapanamjeri et al., (2017). Theorem 3.1 requires a smaller $\alpha$ than in Cherapanamjeri et al., (2017) by a factor of $\mathcal{O}(\kappa)$. Table 1 summarizes the sample complexity and corruption rate requirements of different methods.
@@
Next, for constructing a suitable initialization the requirements on $\alpha$ and $p$ in Theorem 3.2 are more stringent than those for the recovery guarantee of Theorem 3.1. This separation is common in other works on matrix completion (Zheng and Lafferty,, 2016; Yi et al.,, 2016; Zilber and Nadler,, 2022). In terms of sample complexity Theorem 3.2 still improves upon the results in Yi et al., (2016) and Wang and Wei, (2024) and in some parameter regimes also upon the result in Cherapanamjeri et al., (2017). In terms of the fraction of outliers the condition in Theorem 3.2 is more stringent than in Yi et al., (2016) and Cherapanamjeri et al., (2017). Since empirically RGNMR succeeds even with random initialization we believe the condition imposed by Theorem 3.1 can be relaxed. This would allow for less stringent conditions in Theorem 3.2. We leave this for future work.
4 Numerical results
This section details the simulations conducted to compare the various RMC methods. The results are illustrated in Figures 1, 2, 3 and in Appendix B. Each instance consisted of generating a random matrix $L^{*}\in\mathbb{R}^{n_{1}\times n_{2}}$ of a given rank $r$, condition number $\kappa$, subset of observed entries $\Omega\subseteq[n_{1}]\times[n_{2}]$ and a corruption matrix $S^{*}$. We generate $L^{*}$ following Tong et al., (2021). Specifically, we construct $U\in\mathbb{R}^{n_{1}\times r},V\in\mathbb{R}^{n_{2}\times r}$ with entries i.i.d. from the standard normal distribution and orthonormalize their columns. We then construct $\Sigma\in\mathbb{R}^{r\times r}$ as a diagonal matrix with $r$ evenly spaced values between $1$ and $\frac{1}{\kappa}$ on the diagonal. We set $L^{*}=U\Sigma V^{\top}$. Next, to generate $\Omega$ we follow Zilber and Nadler, (2022). Given an oversampling ratio $\rho$ we sample $\Omega$ of size $\lfloor\rho r(n_{1}+n_{2}-r)\rfloor$ randomly without replacement and verify that it contains at least $r$ entries in each column and row. To construct the corruption matrix $S^{*}$ we first sample its support $\Lambda_{*}\subseteq\Omega$ as follows. Given a fraction $\alpha$ of corrupted entries we sample $\lfloor\alpha\cdot|\Omega|\rfloor$ entries out of $\Omega$ randomly without replacement. We then verify that the set $\Omega\setminus\Lambda_{*}$ contains at least $r$ entries in each column and row. Similar to Cai et al., (2024) and Wang and Wei, (2024), we sample the value of each corruption entry uniformly between $\max_{i,j}|L^{*}_{i,j}|$ and $\max_{i,j}|L^{*}_{i,j}|$. Each method is then provided with the matrix $\mathcal{P}_{\Omega}(L^{*}+S^{*})$, the set of entries $\Omega$ and the matrix rank $r$.
We compare RGNMR to AOP (Yan et al.,, 2013), RMC (Cambier and Absil,, 2016), RPCA-GD (Yi et al.,, 2016), RRMC (Cherapanamjeri et al.,, 2017), HUB (Ruppel et al.,, 2020) and HOAT (Wang et al.,, 2023). In addition, as described in Section 2.1, we compare RGNMR to RGNMR-BS. Implementations of all methods except RRMC are publicly available. An implementation of RRMC was kindly provided to us by Wang and Wei, (2024). MATLAB and Python implementations of RGNMR are available at github.com/eilon96/RGNMR. All simulations were run on a 2.1GHz Intel Core i7 CPU with 32GB of memory.
Each algorithm was executed with its default parameters, with the following exceptions: (i) To improve RPCA-GD and HUB performance we increased their number of iterations by using stricter convergence threshold than their default; (ii) In RPCA-GD we employed an over removal factor $\gamma$ of $4$ as it significantly improved its performance in our simulations; (iii) We tuned the parameter $\xi$ in HOAT to $30$; (iv) We provided AOP the exact number of corrupted entries; (v) For RRMC we use the same parameters used in Wang and Wei, (2024).
In Appendix B we illustrate the performance of RGNMR under a broad set of additional scenarios. Specifically, Figure 6 shows that RGNMR can handle a large fraction of corrupted entries. Figures 7 and 8 show that RGNMR outperforms other methods also under non-uniform sampling. Figures 9, 10 and 11 illustrate its performance under a combination of additive noise and outliers. Figure 12 demonstrates that RGNMR can recover matrices with a high rank even if the number of observed entries is relatively small. Figure 13 illustrates that RGNMR performs well on a real dataset involving background extraction in a video. This is a standard benchmark for RMC methods (Yi et al.,, 2016; Cherapanamjeri et al.,, 2017; Huang et al.,, 2021; Cai et al.,, 2024). Finally, Figure 14 shows the runtime of RGNMR and illustrates its performance on large matrices.
5 Limitations and future research
In this section we discuss the theoretical and empirical limitations of RGNMR. These suggest several promising directions for future research.
@@ 5 Limitations and future research
On the theoretical front, though our analysis improves upon existing results, it does not fully account for the remarkable empirical performance of RGNMR. For example, the conditions in Theorem 3.1 on both $\alpha$ and $p$ depend on the condition number $\kappa$ while empirically the performance of RGNMR is not affected by $\kappa$. In addition, our analysis did not consider the case of overparameterization. These are open directions for future theoretical analysis of RGNMR.
On the empirical front, although RGNMR performs well in cases where most RMC methods fail, it is more computationally demanding than some of them. This is because each optimization step of RGNMR requires to solve a large system of linear equations (1), whereas some other methods require only a gradient descent step. As we illustrate in Figure 14, the runtime of RGNMR grows quadratically with $n$. Improving the runtime of RGNMR is therefore an interesting direction for future research.
@@
Acknowledgments
B.N. is the incumbent of the William Petschek Professorial Chair of Mathematics. The research of B.N. was supported in part by grant 2362/22 from the Israel Science Foundation. We thank Tianming Wang and Ke Wei for sharing the code of their work (Wang and Wei,, 2024). We also thank the anonymous reviewers for their valuable feedback that improved the quality of our manuscript.
@@ Appendix A Initialization procedure
  • $\{X_{i,j}\mid(i,j)\in\Omega\}$ - observed entries
@@
  • $r$ - rank of $L^{*}$
  • $p$ - probability of observing an entry
  • $\alpha$ - the maximal fraction of outliers in each row/column
@@
Theorem 3.1 shows that starting from a sufficiently accurate balanced initialization with bounded row norms RGNMR recovers $L^{*}$. Therefore, the goal of our initialization procedure is to construct such initialization. Our proposed scheme starts by constructing an initial estimator of $S^{*}$, which we denote by $S_{init}$. To this end, the function $\mathcal{T}_{\alpha}$ (defined in eq.˜6) is applied to $\mathcal{P}_{\Omega}(X)$. Formally,
@@
$\textnormal{b-SVD}_{r}(\frac{1}{p}(\mathcal{P}_{\Omega}(X)-S_{init}))=\begin{pmatrix}\tilde{U}\tilde{\Sigma}^{\frac{1}{2}}\\ \tilde{V}\tilde{\Sigma}^{\frac{1}{2}}\end{pmatrix}=\begin{pmatrix}U\\ V\end{pmatrix}$. (17)